# NoRA: Nested Low-Rank Adaptation for Efficient Fine-Tuning Large Models

### v1

@@ Abstract
In this paper, we introduce Nested Low-Rank Adaptation (NoRA), a novel approach to parameter-efficient fine-tuning that extends the capabilities of Low-Rank Adaptation (LoRA) techniques. Vanilla LoRA overlooks pre-trained weight inheritance and still requires fine-tuning numerous parameters. To addresses these issues, our NoRA adopts a dual-layer nested structure with Singular Value Decomposition (SVD), effectively leveraging original matrix knowledge while reducing tunable parameters. Specifically, NoRA freezes the outer LoRA weights and utilizes an inner LoRA design, providing enhanced control over model optimization. This approach allows the model to more precisely adapt to specific tasks while maintaining a compact parameter space. By freezing outer LoRA weights and using an inner LoRA design, NoRA enables precise task adaptation with a compact parameter space. Evaluations on tasks including commonsense reasoning with large language models, fine-tuning vision-language models, and subject-driven generation demonstrate NoRA’s superiority over LoRA and its variants. Notably, NoRA reduces fine-tuning parameters|training-time|memory-usage by 4%|22.5%|20.7% compared to LoRA on LLaMA-3 8B, while achieving 2.2% higher performance. Code will be released upon acceptance.
@@ 4.1 Fine-tuning of Large Language Models
Implementation Details. We employed the LLaMA3 8B [45] and LLaMA 7B models, each fine-tuned using a variety of parameter-efficient methods to enhance their commonsense reasoning capabilities. We utilized the Commonsense170K dataset [26] for targeted fine-tuning, which is designed to enhance the models’ understanding of commonsense knowledge across different contexts. The main objective was to assess the effectiveness of each fine-tuning approach in improving the model’s performance on a range of commonsense reasoning tasks.
The baseline models and their respective methodologies for fine-tuning included:
  • LoKr [49]: Utilizes Kronecker products to perform a structured decomposition of the AB matrix, aiming to reduce parameter overhead while maintaining computational efficiency.
  • LoHa [49]: Adapts conventional LoRA by bifurcating it into two distinct branches that implement the Hadamard product, thereby enriching the model’s parameter interaction and adaptation capabilities.
  • FLoRA [12]: Incorporates an additional core via Tucker decomposition to ensure topological consistency with the original model space, thereby facilitating more nuanced adaptation and potentially improving generalization across diverse reasoning tasks.
  • AdaLoRA [54]: Employs singular value decomposition to parameterize incremental updates to the pretrained weight matrices, striking a balance between adaptation fidelity and the preservation of pre-existing knowledge structures.
  • DoRA [33]: Decomposes pretrained weights into their magnitude and directional components, separately fine-tuning each to enhance both the depth and the granularity of the model’s reasoning capabilities.
Post-fine-tuning, model performance was evaluated using a suite of eight benchmark tests focused on commonsense reasoning, including ARC-e, OBQA, SIQA, and more.

### v2

@@ Abstract
In this paper, we introduce Nested Low-Rank Adaptation (NoRA), a novel approach to parameter-efficient fine-tuning that extends the capabilities of Low-Rank Adaptation (LoRA) techniques. Vanilla LoRA overlooks pre-trained weight inheritance and still requires fine-tuning numerous parameters. To addresses these issues, our NoRA adopts a dual-layer nested structure with Singular Value Decomposition (SVD), effectively leveraging original matrix knowledge while reducing tunable parameters. Specifically, NoRA freezes the outer LoRA weights and utilizes an inner LoRA design, providing enhanced control over model optimization. This approach allows the model to more precisely adapt to specific tasks while maintaining a compact parameter space. By freezing outer LoRA weights and using an inner LoRA design, NoRA enables precise task adaptation with a compact parameter space. Evaluations on tasks including commonsense reasoning with large language models, fine-tuning vision-language models, and subject-driven generation demonstrate NoRA’s superiority over LoRA and its variants. Code will be released upon acceptance.
@@ 4.1 Fine-tuning of Large Language Models
Implementation Details. We employed the LLaMA3 8B [45] and LLaMA 7B models, each fine-tuned using a variety of parameter-efficient methods to enhance their commonsense reasoning capabilities. We utilized the Commonsense170K dataset [26] for targeted fine-tuning, which is designed to enhance the models’ understanding of commonsense knowledge across different contexts. The main objective was to assess the effectiveness of each fine-tuning approach in improving the model’s performance on a range of commonsense reasoning tasks. Post-fine-tuning, model performance was evaluated using a suite of eight benchmark tests focused on commonsense reasoning, including ARC-e, OBQA, SIQA, and more.